OpenStudy (anonymous):

double integral question: 1/(x^2+1) R is the triangular region with vertices (0,0), (1,1), and (0,1)

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{x}\frac{ 1 }{ x^2 + 1} dA\]

OpenStudy (anonymous):

I ended up with half of the posted answer = 1/2 ln 2 but it is supposed to be \[\Pi/4 - (1/2 \ln 2)\]

OpenStudy (anonymous):

looks like triple integral and polar coordinates? If anyone has Stewart it is the chapter 12 review question 18 but I don't see how I'm supposed to guess that it is triple integral?